OpenStudy (anonymous):

urn contains 4 white and 6 red rolls. Four balls are drawn at random (without replacement) from the urn. Find the probability distribution of number of white balls?

OpenStudy (ash2326):

Probability distribution W 0 1 2 3 4 P(W) We need to find the probability for each of the case, then we need to fill the table. Do you get this @msingh

OpenStudy (anonymous):

yup

OpenStudy (ash2326):

First case 0, white, It;s easy Can you find the probability of getting all 6

OpenStudy (ash2326):

red

OpenStudy (anonymous):

hmmm. no

OpenStudy (ash2326):

Choosing 4 out of 6 red= 6C4 Choosing 4 out of 10=10C4 \[P(W=0)=\frac{^6C_4}{^{10}C_4}\] Do you get this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

okay when no white balls --- (3/5)^4 = 81/625 one white ball -- C(4,1) (2/5) (3/5)^3 = 216/625 two white balls -- C(4,2) (2/5)^2 (3/5)^2 = 216/625 three white balls -- C(4,3) (2/5)^3 (3/5) = 96/625 four white balls -- (2/5)^4 = 16/625

OpenStudy (anonymous):

is it right

OpenStudy (ash2326):

How did you find 1 white?

OpenStudy (anonymous):

i copy from somewhere

OpenStudy (ash2326):

I think you have replaced them back. Yes copying never helps

OpenStudy (anonymous):

k

OpenStudy (ash2326):

Did you understand how I found probability for first case W=0?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

For second case W=1 R=3 \[P(W=1)=\frac{^4C_1\times^6C_3}{^{10}C_4}\] 1 blue out of 4, 3 red out of 6 4 out of 10

OpenStudy (anonymous):

k

OpenStudy (ash2326):

@msingh are you getting this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@ash2326 where r u